The sum of three numbers is thirty. The second number is twice the first number, and the third number is 5 less than the second number. Find the three numbers.
Answer by nerdybill(7384)

The sum of three numbers is thirty. The second number is twice the first number, and the third number is 5 less than the second number. Find the three numbers.
.
Let x = first number
then from "The second number is twice the first number"
2x = second number
and from "third number is 5 less than the second number"
2x-5 = third number
.
we get our equation from "sum of three numbers is thirty"
x + 2x + 2x-5 = 30
5x-5 = 30
5x = 35
x = 7 (first number)
.
second number:
2x = 2(7) = 14
.
third number:
2x-5 = 2(7)-5 = 14-5 = 9